# Relative energy constraint in Boltzmann’s surface entropy eliminates thermodynamic inconsistencies and allows negative absolute temperatures

### v1

@@ Abstract
Quantum harmonic oscillator. For a quantum harmonic oscillator, the energy of the system is given as [40] $E_{n}=h\nu\left(n+\frac{1}{2}\right)$, where $n=0,1,2,...$ and $\nu$ represents the frequency of vibration of the oscillator. It follows that $\Omega=1+n=\frac{E}{h\nu}+\frac{1}{2}$. At the high-temperature limit, the states can be treated as a continuum, so that $k_{B}T_{G}=\frac{h\nu}{2}+E$. Although applying the original definition of Boltzmann entropy leads to infinite temperature, the newly proposed definition (Eq. (5)) results in $\omega_{r}=\frac{\epsilon_{r}E}{h\nu}$, leading to $E=k_{B}T_{B_{r}}$, which predicts finite, positive temperatures. Again, only the relative-energy-constrained Boltzmann entropy leads to predictions consistent with the canonical ensemble (Table 1)! One can also derive expressions valid for the temperature under all cases, using discretized definitions of the partition function and temperature (see, e.g., ref. [41]), as shown in the Supplementary Information [42] and in Table 1. The ensuing $E$ vs. $T$ relationships for various entropy choices are plotted in Figure 2. Therein, one sees that only the modified Boltzmann entropy predicts thermodynamically consistent results at all temperatures, including the existence of the zero-point energy $\left(\frac{1}{2}h\nu\right)$ and agreeing with the high-temperature prediction of $E=k_{B}T$. On the contrary, while the Gibbs entropy fails to predict the existence of a zero-point energy (i.e., leads to a Planck oscillator rather than a Schrodinger oscillator [14]!), the Boltzmann entropy predicts infinite temperatures for all energies. Note that, to the best of the author’s knowledge, this is the first derivation of the temperature of a single, isolated quantum harmonic oscillator, i.e., in the microcanonical ensemble; previous expressions relied on the consideration of multiple oscillators, in the limit of a large number of oscillators (see, e.g., [43] and [44]).
@@
Figure 2: Comparison of various microcanonical entropies in terms of their ability to predict the thermodynamic behavior of a quantum harmonic oscillator. The canonical prediction is shown using a blue line and the high-temperature canonical/modified Boltzmann predictions are shown using a dashed black line. Blue, red, and green circles represent predictions of the modified Boltzmann, Gibbs, and conventional Boltzmann entropies in the microcanonical ensemble. In these cases, lines are simply a guide. Only the modified Boltzmann entropy predicts the existence of the correct zero-point energy at 0 K and degenerates to the limit of $k_{B}T$ at high temperatures.
@@
Quantum particle in a 1D box. The energy of a quantum particle in a one-dimensional (1D) box is given as [40] $E_{n}=\frac{an^{2}}{L^{2}}$, where $a=\frac{\hbar^{2}\pi^{2}}{2m}$ and $n=1,2,3,...$, so that $\Omega=L\sqrt{\frac{E}{a}}$. At high temperatures, the states form a continuum, and we find that $E=\frac{k_{B}T_{G}}{2}$ and $P_{G}=T_{G}\frac{\partial S_{G}}{\partial L}=\frac{2E}{L}$. This statistical-mechanical pressure equals the thermodynamic pressure calculated as $P_{T}=-\frac{\partial E}{\partial L}$. Now, although the conventional Boltzmann entropy leads to a negative temperature $\left(E=-\frac{k_{B}T_{B}}{2}\right)$ and negative pressure $\left(P_{B}=-\frac{2E}{L}\right)$, the newly proposed definition of Boltzmann entropy fixes these predictions, as seen in Table 1, to yield $E=\frac{k_{B}T_{B_{r}}}{2}$ and $P_{B_{r}}=\frac{2E}{L}$.
@@
Two-level system and the possibility of negative temperatures. We now demonstrate that the newly proposed definition of Boltzmann entropy still allows for negative temperatures for systems having non-monotonic density of states [21, 22, 12], i.e., a bounded phase space. Considering the well-known case of a two-state system, with energy levels 0 and $E_{1}$ and $N_{0}$ particles occupying the ground state and $N_{1}$ particles occupying the excited state, the energy of the system is $E=N_{1}E_{1}$ and the number of particles is $N=N_{0}+N_{1}$. It follows that the number of states with energy less than or equal to $E$ is
$\Omega=\sum_{N_{0}^{\prime}=1}^{N_{1}}\frac{N!}{N_{0}^{\prime}!N_{1}^{\prime}!}$
where $N_{1}=E/E_{1}$ and $N_{1}^{\prime}=N-N_{0}^{\prime}$. The original Boltzmann partition function is given as
@@
where $\Delta\Omega$ is obtained as the difference in the summation expressions of $\Omega$ with $N_{0}^{\prime}$ going up to $(N_{1}+1)$ and $N_{1}$. Considering a large number of particles and applying Stirling’s approximation results in
$S_{B}=k_{B}N\ln N-k_{B}(N_{0}-1)\ln(N_{0}-1)-$
$k_{B}(N_{1}+1)\ln(N_{1}+1)+k_{B}\ln\frac{\epsilon}{E_{1}}.\;\;\;\;\;\;\;$
@@
$T_{B}=\frac{\Delta E}{\Delta S_{B}}=\frac{E_{1}}{k_{B}\ln\left(\frac{(N_{0}-1)^{N_{0}-1}(N_{1}+1)^{N_{1}+1}}{(N_{0}-2)^{N_{0}-2}(N_{1}+2)^{N_{1}+2}}\right)}.$
Since we are considering a large number of particles, the temperature reduces to $T_{B}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$. Now, if $E_{1}>0$ and $N_{0}<N_{1}$, i.e., we consider a system exhibiting “population inversion”, we find that $T_{B}<0$. Analogously, the modified Boltzmann partition function is given as
@@
This leads to, via the Stirling’s approximation, the following expression for the temperature:
$T_{B_{r}}=\frac{\Delta E}{\Delta S_{B_{r}}}=\frac{E_{1}}{k_{B}\ln\left(\frac{(N_{0}-1)^{N_{0}-1}(N_{1}+1)^{N_{1}+2}}{(N_{0}-2)^{N_{0}-2}(N_{1}+2)^{N_{1}+2}N_{1}}\right)},$
@@
which also simplifies, for a large number of particles, to $T_{B_{r}}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$, also allowing negative temperatures. (This is the same expression that one obtains by direct enumeration as well [12].) On the other hand, Gibbs entropy would not admit negative temperatures at all due to a monotonically increasing total number of states. Thus, Gibbs entropy cannot account for systems with bounded spectra in a thermodynamically consistent manner [45, 46, 12]. Note that this also indicates that Gibbs and Boltzmann entropies can disagree, even in the thermodynamic limit. Now, according to the above discussion, the newly proposed relative-energy-constrained Boltzmann entropy ($S_{B_{r}}$) allows for both negative and positive temperatures, and in the appropriate contexts in which they should be seen, i.e., systems with bounded and unbounded phase space, respectively (see Table 1). Therefore, this work indicates that previous negative temperature measurements for various systems are indeed correct [47, 48, 49, 50, 51, 52, 53], unlike what is stated in ref. [5]. In fact, another argument against negative temperature, viz., the incorrect interpretation of Carnot efficiencies being more than one has also been addressed by Abraham and Penrose [54].
@@
Conclusions. In this work, we introduced a relative energy constraint in the calculation of the microcanonical partition function, which eliminates negative/infinite temperatures for systems with unbounded spectra, such as ideal gases with a few degrees of freedom, a quantum harmonic oscillator, and a quantum particle confined in an infinite well. At the same time, the proposed entropy definition allows for negative temperatures for systems with non-monotonic density of states, i.e., a bounded spectrum, such as a two-state system with population inversion. This makes the proposed entropy definition the one that is the closest to predictions from the canonical ensemble in each case considered (see Table 1). The physical insight underlying the proposed approach is that systems with higher energy will interact more with their surroundings, thus causing a larger perturbation in their energy. Mathematically, while an absolute energy constraint uses up one degree of freedom from the system (by fixing the value of $E$), the use of a relative energy constraint fixes only the ratio $H/E$, thus retaining the original number of degrees of freedom of the system and resulting in the correct scaling of energy with temperature for various systems. Thereby, we proposed a resolution to the ongoing debate in microcanonical statistical mechanics on the use of the Gibbs volume entropy and Boltzmann surface entropy. Overall, this work sets the foundation for using the relative-energy-constrained Boltzmann surface entropy in the microcanonical ensemble of statistical mechanics. Thereby, the work puts on firm theoretical footing the existence of negative temperatures for systems with bounded energy spectra. Finally, this work lays a solid foundation for accurate theoretical descriptions of isolated systems with few degrees of freedom using statistical mechanics.
Acknowledgements.
@@
A.G.R. acknowledges the Infosys Foundation, Bengaluru for an Infosys Young Investigator grant. A.G.R. thanks Profs. B. J. Cherayil, K. G. Ayappa, N. M. Dixit, P. K. Maiti, P. R. Nott, and S. N. Punnathanam for insightful discussions, and Profs. A. Gupta and S. K. Gupta for valuable suggestions regarding the manuscript.

### v2

@@ Abstract
Next, for a quantum harmonic oscillator, the energy of the system is given as 52 $E_{n}=h\nu\left(n+\frac{1}{2}\right)$, where $n=0,1,2,...$ and $\nu$ represents the frequency of vibration of the oscillator. It follows that $\Omega=1+n=\frac{E}{h\nu}+\frac{1}{2}$. Since $0\leq n<\infty$, the system’s energy is unbounded. At the high-temperature limit, the states can be treated as a continuum, so that $k_{B}T_{G}=\frac{h\nu}{2}+E$. Although applying the original definition of Boltzmann entropy leads to infinite temperature since $\omega=\frac{\epsilon}{h\nu}$, the newly proposed definition (Eq. (9)) results in $\omega_{r}=\frac{\epsilon_{r}E}{h\nu}$, leading to $E=k_{B}T_{B_{r}}$, which predicts finite, positive temperatures. Again, only the relative-energy-constrained Boltzmann entropy leads to predictions consistent with the canonical ensemble (Table 1)! Note that the degeneracy entropy is zero for every energy level because $\Delta\Omega=1$, indicating that the degeneracy entropy would predict $T_{D}\rightarrow\infty$, where $T_{D}$ denotes the temperature corresponding to the degeneracy entropy, similar to the prediction of the Boltzmann entropy in this case.
One can also derive expressions valid for the temperature under all cases, using discretized definitions of the partition function and temperature (see, e.g., ref.  53), as shown in the Supporting Information and in Table 1. The ensuing $E$ vs. $T$ relationships for various entropy choices are plotted in Figure 3. Therein, one sees that only the relative-energy-constrained Boltzmann entropy predicts thermodynamically consistent results at all temperatures, including the existence of the zero-point energy $\left(\frac{1}{2}h\nu\right)$ and agreeing with the high-temperature prediction of $E=k_{B}T$. On the contrary, while the Gibbs entropy fails to predict the existence of a zero-point energy (i.e., leads to a Planck oscillator rather than a Schrodinger oscillator; see Figure 3 and ref.  15 for a definition of Planck and Schrodinger oscillators), the Boltzmann entropy predicts infinite temperatures for all energies. Note that, to the best of the author’s knowledge, this is the first derivation of the temperature of a single, isolated quantum harmonic oscillator, i.e., in the microcanonical ensemble; previous expressions relied on the consideration of multiple oscillators, in the limit of a large number of oscillators (see, e.g., refs. 54 and 55). Note that, as seen in Figure 3, in the microcanonical ensemble, only discrete values of $E$, i.e., $E_{0}+\frac{1}{2}h\nu$ (where $E_{0}$ is the zero-point energy) are allowed (as per the spectrum of the harmonic oscillator) as indicated by the blue circles, whereas, in the canonical ensemble, all possible values of $E$ are allowed due to the contact with a thermal reservoir, as depicted by the continuous blue line.
@@
Figure 3: Comparison of various microcanonical entropies in terms of their ability to predict the thermodynamic behavior of a quantum harmonic oscillator. Blue, red, and green circles represent predictions of the proposed, Gibbs, and Boltzmann/degeneracy entropies in the microcanonical ensemble. The canonical prediction is shown using a blue line and the high-temperature canonical prediction is shown using a dashed black line. In these cases, lines are simply a guide. Only the proposed entropy predicts the existence of the correct zero-point energy at 0 K and degenerates to the limit of $k_{B}T$ at high temperatures.
@@
Now, we consider a quantum particle in a one-dimensional (1D) box. The energy of this system is given as 52 $E_{n}=\frac{an^{2}}{L^{2}}$, where $a=\frac{\hbar^{2}\pi^{2}}{2m}$ and $n=1,2,3,...$, so that $\Omega=L\sqrt{\frac{E}{a}}$ and consequently $\omega=\frac{\partial\Omega}{\partial E}\epsilon=\frac{L\epsilon}{2\sqrt{Ea}}$. Since there is no limit on how large $n$ can be, the system has an unbounded energy spectrum. At high temperatures, the states form a continuum, and we find that $E=\frac{k_{B}T_{G}}{2}$ and $P_{G}=T_{G}\frac{\partial S_{G}}{\partial L}=\frac{2E}{L}$. This statistical-mechanical pressure equals the thermodynamic pressure calculated as $P_{T}=-\frac{\partial E}{\partial L}$. Now, although the conventional Boltzmann entropy leads to a negative temperature $\left(E=-\frac{k_{B}T_{B}}{2}\right)$ and negative pressure $\left(P_{B}=-\frac{2E}{L}\right)$, the newly proposed definition of Boltzmann entropy fixes these predictions to yield $\omega_{r}=\frac{L\epsilon_{r}}{2}\sqrt{\frac{E}{a}}$, $P_{B_{r}}=\frac{2E}{L}$, and $E=\frac{k_{B}T_{B_{r}}}{2}$, as seen in Table 1. Again, because $\Delta\Omega=1$, the degeneracy entropy predicts infinite temperature for every energy level (i.e., $T_{D}\rightarrow\infty$), thus not being useful as a definition of entropy.
@@
We now demonstrate that the newly proposed definition of Boltzmann entropy still allows for negative temperatures for systems having non-monotonic density of states 22, 23, 13, i.e., a bounded phase space. Considering the well-known case of a two-level system, with energy levels 0 and $E_{1}$ and $N_{0}$ particles occupying the ground state and $N_{1}$ particles occupying the excited state, the energy of the system is $E=N_{1}E_{1}$ and the number of particles is $N=N_{0}+N_{1}$. This system has an upper bound on its energy equal to $NE_{1}$. It follows that the number of states with energy less than or equal to $E$ is
$\Omega=\sum_{N_{0}^{\prime}=1}^{N_{1}}\frac{N!}{N_{0}^{\prime}!N_{1}^{\prime}!}$ (17)
where $N_{1}=E/E_{1}$ and $N_{1}^{\prime}=N-N_{0}^{\prime}$. We use the discrete definitions of the partition function and temperature so that the derivation holds for both low and high temperatures. Accordingly, the Gibbs temperature is obtained as
$T_{G}=\frac{\Delta E}{\Delta S_{G}}=\frac{E_{1}}{k_{B}\ln\left(\frac{\sum_{N_{0}^{\prime}=1}^{N_{1}+1}\frac{N!}{N_{0}^{\prime}!N_{1}^{\prime}!}}{\sum_{N_{0}^{\prime}=1}^{N_{1}}\frac{N!}{N_{0}^{\prime}!N_{1}^{\prime}!}}\right)}.$ (18)
Moving further, the original Boltzmann partition function is given as
@@
where $\Delta\Omega$ is obtained as the difference in the summation expressions of $\Omega$ with $N_{0}^{\prime}$ going up to $(N_{1}+1)$ and $N_{1}$. Thus
$S_{B}=k_{B}\ln\left(\frac{N!}{(N_{0}-1)!(N_{1}+1)!}\frac{\epsilon}{E_{1}}\right).$
@@
$T_{B}=\frac{\Delta E}{\Delta S_{B}}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}-1}{N_{1}+2}\right)}.$ (19)
For a large number of particles, the temperature can simply be written as $T_{B}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$. Now, if $E_{1}>0$ and $N_{0}<N_{1}$, i.e., we consider a system exhibiting “population inversion”, we find that $T_{B}<0$. Analogously, the relative-energy-constrained Boltzmann partition function is given as
@@
This leads to the following expression for the temperature:
$T_{B_{r}}=\frac{\Delta E}{\Delta S_{B_{r}}}=\frac{E_{1}}{k_{B}\ln\left(\frac{(N_{1}+1)(N_{0}-1)}{N_{1}(N_{1}+2)}\right)},$ (20)
which also simplifies, for a large number of particles, to $T_{B_{r}}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$, also allowing negative temperatures. Note that the degeneracy entropy predicts $T=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}+1}\right)}$ as the number of degenerate states changes from $\frac{N!}{N_{0}!N_{1}!}$ to $\frac{N!}{(N_{0}-1)!(N_{1}+1)!}$ as the energy changes from $N_{1}E_{1}$ to $(N_{1}+1)E_{1}$.
Figure 4 shows the temperature predictions made by the proposed, Boltzmann, Gibbs, degeneracy, and canonical (see the Supporting Information) entropies, for a two-level system consisting of $N=170$ particles, where the predictions are obtained numerically. Although the factorials disappear in the temperature expressions for entropies other than the Gibbs entropy, as seen above, the factorials are retained in the Gibbs entropy expression. The chosen $N=170$ is thus the largest one for which $N!$ is calculable in double-precision floating-point format. In Figure 4, one sees that the proposed entropy leads to a slightly lower temperature than the conventional Boltzmann entropy. Moreover, we see that Gibbs entropy does not admit negative temperatures at all due to a monotonically increasing total number of states. Instead, for system configurations wherein $T_{B}$ and $T_{B_{r}}$ are negative, $T_{G}$ is exceptionally large, which actually represents a cooler system since negative absolute temperatures are hotter than positive absolute temperatures.
Interestingly, while the predictions of the degeneracy entropy are closer to the original Boltzmann entropy, the predictions of the proposed entropy are closer to canonical entropy, considering that the entropy definitions are staggered due to the use of finite differences. These findings support the earlier claims that Gibbs entropy does not account for systems with bounded spectra in a thermodynamically consistent manner 56, 57, 13. They also indicate that Gibbs and Boltzmann entropies can disagree, even in the thermodynamic limit. Moreover, these results provide an experimental route to test the validity of the proposed and Boltzmann entropies by measuring the temperature of a two-level system.
@@
Note that, according to the above discussion, the newly proposed relative-energy-constrained Boltzmann entropy ($S_{B_{r}}$) allows for both negative and positive temperatures, and in the appropriate contexts in which they should be seen, i.e., systems with bounded and unbounded phase space, respectively (see Table 1). Therefore, this work indicates that previous negative temperature measurements for various systems are indeed correct 58, 59, 60, 61, 62, 63, 64, unlike what is stated in ref.  5. In fact, another argument against negative temperature, viz., the incorrect interpretation of Carnot efficiencies being more than one has also been addressed by Abraham and Penrose 65. The reader is also referred to the comprehensive review by Baldovin et al., which lays out the statistical mechanics of systems with negative absolute temperatures and discusses how they are consistent with both equilibrium and non-equilibrium thermodynamics.66
Figure 4: Comparison of the predictions made by various microcanonical entropies for a two-level system consisting of $N=170$ particles. $E_{1}$ and $N_{1}$ denote, respectively, the energy and the number of particles in the excited state, and $N_{0}$ denotes the number of particles in the ground state, as shown in the bottom-left inset. The top-right inset shows the Gibbs entropy’s temperature prediction on a log scale.
@@
To conclude, we introduced a relative energy constraint in the calculation of the microcanonical partition function, which eliminates negative/infinite temperatures for systems with unbounded spectra (i.e., without an upper limit on the allowed energy), such as ideal gases with a few degrees of freedom, a quantum harmonic oscillator, and a quantum particle confined in an infinite well. At the same time, the proposed entropy definition allows for negative temperatures for systems with non-monotonic density of states, i.e., a bounded spectrum (with an upper limit on energy), such as a two-level system with population inversion. This makes the proposed entropy definition the one that is the closest to predictions from the canonical ensemble in each case considered (see Table 1), although such equivalence was not imposed in this work and may not arise for all possible systems, as shown for the two-level system. We also explained how the degeneracy-based entropy is not applicable in many cases and also does not agree with canonical-ensemble predictions.
The physical insight underlying the proposed relative-energy-constrained entropy is that systems with higher energy will interact more with their surroundings, thus causing a larger perturbation in their energy. This proposal is motivated using the fact that the interactions of a system with its surroundings can never be fully eliminated, combined with the Heisenberg uncertainty principle and an eigenstate thermalization time inversely proportional to the system energy. Mathematically, while an absolute energy constraint uses up one degree of freedom from the system (by fixing the value of $E$), the use of a relative energy constraint fixes only the ratio $H/E$, thus retaining the original number of degrees of freedom of the system and resulting in the correct scaling of energy with temperature for various systems. We show that this is critical to describing correct physics, e.g., recovering the zero-point energy of a quantum harmonic oscillator in the microcanonical ensemble and avoiding unphysical temperatures. Nevertheless, we remark that, depending on the underlying physics of the system at hand, one would have to carefully make a choice between applying an absolute or a relative energy constraint. Another open question is determining the value of $\epsilon_{r}$ for a given microcanonical system very weakly coupled to a bath.
In summary, we proposed a resolution to the ongoing debate in microcanonical statistical mechanics on the use of the Gibbs volume entropy versus Boltzmann surface entropy. In this regard, we also proposed an experimental route to test the validity of the proposed and the Boltzmann entropies using a two-level system. Overall, this work sets the foundation for using the relative-energy-constrained Boltzmann surface entropy in the microcanonical ensemble of statistical mechanics. Therefore, the work puts on a firm theoretical footing the existence of negative temperatures for systems with bounded energy spectra. Finally, this work lays a solid foundation for accurate theoretical descriptions of isolated systems with few degrees of freedom using statistical mechanics.
{acknowledgement}
@@
A.G.R. acknowledges the Infosys Foundation, Bengaluru for an Infosys Young Investigator grant. A.G.R. thanks Profs. Subroto Mukerjee, Sumilan Banerjee, Sudeep N. Punnathanam, Binny J. Cherayil, K. Ganapathy Ayappa, Narendra M. Dixit, Prabal K. Maiti, and Prabhu R. Nott, and Dr. Garrett R. Dowdy for insightful discussions, and Profs. Diptiman Sen, Ankur Gupta, and Sanjeev K. Gupta for valuable comments/suggestions regarding the manuscript.
{suppinfo}
Derivation of the energy expression for a quantum harmonic oscillator in the microcanonical ensemble valid at all temperatures, high-temperature limits for the energy of a quantum harmonic oscillator, examination of the energy-dependent energy window proposed in previous work, the two-level system in the canonical ensemble, and examining the consistency condition mentioned in previous work.